# Localization behavior in a Hermitian and non-Hermitian Raman lattice

### v1

@@ Abstract
In most localization behavior studies, a crucial consideration lies in the necessity for the initial state to populate only a few lattice sites to ensure clarity in distinguishing between different phases (Fig.3(c)). While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifierMurthy et al. (2014); Asteria et al. (2021) or a quantum gas microscope Bakr et al. (2009); Sherson et al. (2010); Miranda et al. (2015); Haller et al. (2015); Edge et al. (2015); Omran et al. (2015); Cheuk et al. (2016); Parsons et al. (2016); Yamamoto et al. (2016); Mitra et al. (2018); Yamamoto et al. (2020); Kwon et al. (2022), this can introduce complexity and inconvenience to the experimental arrangement. An intriguing observation emerges from numerical simulations represented in Fig. 3(a), indicating that in the extended phase, the presence of spin-down atoms is negligible when all initial states are occupied by spin-up atoms. Contrastingly, in the critical phase, clear evolution of spin-down atoms is observable over time. This can provide another way to detect the emergence of a critical phase in the current system.
In Fig.4(a), the spin polarization for systems where all lattice sites are initially uniformly occupied by spin up atoms is depicted for various phases. Within the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while in critical and localized states, distinct oscillations occur, with the critical phase displaying larger oscillation amplitudes than the localized phase. Comparing with the distinct boundary between the extended and critical phases, the demarcation between the critical and localized phases, while not as pronounced, remains identifiable. Utilizing time-averaged spin polarization can aid in characterizing the emergence of the critical phase (Fig.3(c)-(d)). Notably, the time-averaged spin polarization tends towards zero in the critical phase, contrasting with higher values outside this phase.
@@
Figure 4: Evolution of spin polarization for different phases. (a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$) with the initial state comprising spin-up atoms in a uniform distribution. (b) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state uniform distribution. (c) Time-averaged spin polarization as a function of $M_{z}$ averaged over $t=0-200\text{s}$, which show minimal disparity between initial states of uniform distribution or Gaussian distribution. (d) Time-averaged spin polarization as a function of $M_{z}$ and $t_{so}$. All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
IV IV. Coexistence of different states
@@
Returning to the spin-independent incommensurate lattice case, although the critical phase is absent in such a system, under specific conditions, extended and localized states can coexist, separated by a mobility edge. For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions Wang et al. (2022). Illustrated in Fig.5(a) is the fractal dimension corresponding to various eigenenergies and $M_{z}$, revealing regions where three or two distinct types of eigenstates coexist, separated by mobility edges. To validate the coexistence of difference phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.5(b)). Notably, as the system size increases, $\eta$ converges towards 0 for all states in zones I and IV, indicative of localization states, while tending towards 1 for all states in zone III, signifying extended states. Conversely, within zones II and V, the fractal dimension markedly deviates from 0 and 1 and remains relatively constant with system size, suggesting a critical state in these regions.
@@
The time evolution of the wave packet under such circumstances, as depicted in Fig.5(c), showcases characteristics signatures of these three distinct phases concurrently. On the whole, the time evolution of the wave packet exhibits quasi-localization, highlighting the presence of critical states. However, the higher concentration of atom clouds in the vicinity of central sites, particularly for spin-down atoms, indicates the existence of localized states. Furthermore, the inset of Fig.5(c) illustrates instances of ballistic expansion behavior of localized states upon closer inspection of the color bar. Through simulations of time evolution, the coexistence of extended, critical, and localized states under such conditions is further substantiated.
@@
Figure 5: Phase with coexisting localized, extended, and critical zones. (a) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$ with $t_{so}=0.3$. Simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition. (b) Fractal dimension of different eigenstates as a function of eigenenregy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. (c) Time evolution of the wave packet for the system located in the phase with coexisting localized, extended, and critical zones ($M_{z}=1.5$, $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$). Simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
V V. Suppression of critical phase in non-Hermitian Raman lattice
@@
Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model Jiang et al. (2019) and the PT-symmetric Aubry-André model Yuce (2014); Liang et al. (2014); Hang et al. (2015); Harter et al. (2016); Zeng et al. (2017); Longhi (2019). Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum Longhi (2014); Yuce (2014); Liang et al. (2014); Harter et al. (2016), topological phase transitions Longhi (2019), and the localization characteristics of eigenstates Hatano and Nelson (1996); Shnerb and Nelson (1998); Hang et al. (2015); Zeng et al. (2017); Jiang et al. (2019); Zhou (2023). Given these insights, it becomes pertinent to investigate the effects of non-Hermiticity on our system.
@@
Expanding our setup into the non-Hermitian regime involves introducing dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam Zhou (2023). We focus on the scenario with $M_{z,\uparrow}/M_{z,\downarrow}=-1$, which exhibits a critical phase in the Hermitian context. The phase diagram for the system with dissipation ($\gamma_{\downarrow}=-2$) is illustrated in Fig. 6(a), revealing a transition of critical phases in the Hermitian regime to extended or localized phases in non-Hermitian regime. Fig. 6(b) presents the mean fractal dimension $\bar{\eta}$ for varying dissipation values $\gamma_{\downarrow}$ with $t_{so}={1.0}$. The plateau characterizing the critical phase in the Hermitian setup vanishes upon the introduction of dissipation. For system with larger $M_{z}$, the mean fractal dimension decreases, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises, signifying a transition towards extended states.
Furthermore, Fig. 6(c) shows the fraction of eigenstates exhibiting different fractal dimensions, depicting a critical-extended transition. Employing a methodology akin to that in Fig. 5, it is observed that in the Hermitian regime, the fractal dimension remains stable, indicative of critical states in these regions. However, upon introducing dissipation ($\gamma_{\downarrow}=2$), $\eta$ converges towards 1 for all states when the system size increases, confirming the transformation of nearly all states into extended states and the suppression of the critical phase within the system due to the induced dissipation (Fig. 6(d)). This suppression of the critical phase is attributed to non-Hermiticity, as evidenced by the persistence of this effect even under conditions where $\gamma_{\uparrow}/\gamma_{\downarrow}=-1$.
@@
Figure 6: Suppression of critical phase in non-Hermitian Raman lattice. (a) The phases diagram after inducing dissipation ($\gamma_{\downarrow}=-2$) to the system, distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. The dashed lines show the boundary between different phases in Hermitian regime. (b) Mean fractal dimension $\bar{\eta}$ for different dissipation $\gamma_{\downarrow}$ as a function of $M_{z}$ with $t_{so}={1.0}$. (c) Fraction of eigenstates with $\eta>0.8$, $\eta<0.3$ and $0.80.3<\eta<$ 0.8 respectively as a function of $\gamma_{\downarrow}$ with $t_{so}=1.0$ and $M_{z}=1.0$. (d) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. The top one is without dissipation and the bottom one is with dissipation. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. Unless otherwise specified, all simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition.
VI VI. Summary
@@
This study introduces a versatile Raman lattice configuration aimed at investigating the localization behavior for alkaline-earth-like atoms, exemplified by 173Yb. The experimental setup explores three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime. In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, the emergence of a critical phase is observed, which can be detected by expansion dynamics and spin polarization evolution. Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, a new phase with coexistence of extended, critical, and localized phases separated by mobility edges shows up. On the other hand, we also find that introducing dissipation into the system leads to the suppression of the critical phase due to non-Hermiticity within specific parameter spaces. This work aims to stimulate further experimental investigations into the interplay between the localization behavior and non-Hermiticity Zhou (2023); Jiang et al. (2019) in ultracold atoms.
VII ACKNOWLEDGMENTS
@@
GBJ acknowledges support from the RGC through 16306922, 16302123, 16305024, N-HKUST636-22, and RFS2122-6S04.

### v2

@@ IV Spin dynamics and density imbalance
In most localization behavior studies, a crucial consideration lies in the necessity for the initial state to populate only a few lattice sites to ensure clarity in distinguishing between different phases (Fig.3(c)). While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifier [53, 6] or a quantum gas microscope [8, 62, 47, 22, 16, 55, 14, 56, 75, 49, 76, 32], this can introduce complexity and inconvenience to the experimental arrangement. To address this constraint, we also numerically simulate the evolution of spin dynamics and density imbalance across varied initial conditions, thereby offering complementary pathways to distinguish phases within the current system.
We begin by examining the initial state in which all lattice sites are occupied by spin-up atoms ($|{\psi}\rangle=|{\uparrow\uparrow\uparrow\cdots\uparrow}\rangle$). In the single-particle picture, this state corresponds to a momentum distribution localized at $k=0$. Figure 4(a), presents the time evolution of the spin polarization $I_{p}=(n_{\uparrow}-n_{\downarrow})/(n_{\uparrow}+n_{\downarrow})$ under this configuration. In the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while pronounced oscillations are observed in both the critical and localized phases. Utilizing time-averaged spin polarization can aid in characterizing the transition between the critical phase and extend phase (Fig.4 (b)). Specifically, it approaches zero in the critical phase but remains close to one in the extended phase. To elucidate the origin of the distinct spin dynamics across the extended–critical transition, we simulate the reconstructed momentum distribution during the evolution (Fig.4(c)–(d)). In the extended phase, the momentum distribution remains localized near $k=0$ throughout the evolution. By contrast, in the critical phase, the distribution broadens to momentum far from $k=0$. In the absence of a quasi-periodic lattice, the system is described by the Bloch Hamiltonian [82, 65, 87, 41]:
$\mathcal{H}(k)=(m_{z}-2t_{0}\cos{k})\sigma_{z}+2t_{\text{so}}\sin{k}~\sigma_{y}$ (7)
At $k=0$, the off-diagonal term vanishes, leading to the suppression of spin dynamics at this momentum.
Notably, in the present system, the transition between the critical and localized phases coincides with the topological phase boundary [69]. Away from the topological regime, the system exhibits detuned spin oscillations. Although distinctions in spin dynamics do exist between the critical and localized phases, these differences are less pronounced than the clear demarcation observed between the extended and critical phases. Consequently, spin dynamics may not serve as a reliable indicator for identifying the critical‑to‑localized phase transition.
@@
Figure 4: Evolution of spin polarization for different phases. (a) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state with uniform distribution. (b) Time-averaged spin polarization as a function of $M_{z}$ and $t_{\text{so}}$. (c, d) Corresponding reconstructed momentum distribution during the time evolution at extended phase ($M_{z}=0.5$) and critical phase ($M_{z}=2.0$). All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
To further probe the critical‑to‑localized phase transition, we examine an alternative initial state in which spin-up atoms half-fill the lattice, described by $|{\psi}\rangle=|{0\uparrow 0\uparrow 0\uparrow\cdots}\rangle$. This state can be experimentally prepared and detected using an optical superlattice  [61, 45]. Figure 5 (a) and (b) display the time evolution of the density imbalance between even and odd sites, defined as $I_{d,\sigma}=(N_{e,\sigma}-N_{o,\sigma})/(N_{e}+N_{o})$, where $\sigma=\uparrow,\downarrow$ and $N_{e}(N_{o})$ denotes the atom number on even (odd) sites, in both the critical and localized regimes. After long-time evolution, the density imbalance for spin-up atoms remains nonzero in the localized regime, whereas it decays to zero with oscillation in the critical regime. Interestingly, spin-down atoms tend to occupy lattice sites of the opposite parity, as indicated by the opposite sign of their density imbalance. This behavior may originate from the fact that the spin-flip term in our Hamiltonian includes only nearest-neighbor couplings without on-site spin flipping. We can further define the overall density imbalance as $I_{d}=I_{d,\uparrow}-I_{d,\downarrow}$ and show its time average $\bar{I}_{d}$ over the interval t=0200t=0-200 in Fig. 5 (c). A clear boundary emerges between the localized and critical phases. Combined with the time-averaged spin polarization $\bar{I}_{p}$ introduced earlier, both phase transitions can be clearly identified (Fig. 5), even in systems lacking the ability to prepare or detect Gaussian wavepackets localized to only a few lattice sites.
Figure 5: Evolution of density imbalance for different phases. (a, b) Time evolution of the density imbalance for spin-up (a) and spin-down (b) atoms in the critical phase ($M_{z}=2.0,t_{\text{so}}=0.3$) and the localized phase ($M_{z}=5.0,t_{\text{so}}=0.3$). (c) Time-averaged density imbalance as a function of $M_{z}$ and $t_{\text{so}}$. The markers indicate the parameter sets corresponding to the dynamics shown in (a) and (b). (d) Time-averaged spin polarization $\bar{I}_{p}$ and density imbalance $\bar{I}_{d}$ as a function of $M_{z}$ averaged over the time interval t=0200t=0-200. The expected phase transition position is marked by a vertical dashed line. All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
V Coexistence of different states
@@
Returning to the spin-independent incommensurate lattice case, although the critical phase is absent in such a system, under specific conditions, extended and localized states can coexist, separated by a mobility edge. For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions [70]. Illustrated in Fig.6(a) is the fractal dimension corresponding to various eigenenergies and $M_{z}$, revealing regions where three or two distinct types of eigenstates coexist, separated by mobility edges. To validate the coexistence of different phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.6(b)). Notably, as the system size increases, $\eta$ converges towards 0 for all states in zones I and IV, indicative of localization states, while tending towards 1 for all states in zone III, signifying extended states. Conversely, within zones II and V, the fractal dimension markedly deviates from 0 and 1 and remains relatively constant with system size, suggesting a critical state in these regions.
@@
The time evolution of the wave packet under such circumstances, as depicted in Fig.6(c), showcases characteristic signatures of these three distinct phases concurrently. On the whole, the time evolution of the wave packet exhibits quasi-localization, highlighting the presence of critical states. However, the higher concentration of atom clouds in the vicinity of central sites, particularly for spin-down atoms, indicates the existence of localized states. Furthermore, the inset of Fig.6(c) illustrates instances of ballistic expansion behavior of localized states upon closer inspection of the color bar. Through simulations of time evolution, the coexistence of extended, critical, and localized states under such conditions is further substantiated.
@@
Figure 6: Phase with coexisting localized, extended, and critical zones. (a) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$ with $t_{so}=0.3$. Simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition. (b) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. (c) Time evolution of the wave packet for the system located in the phase with coexisting localized, extended, and critical zones ($M_{z}=1.5$, $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$). Simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
VI Localization behavior in non-Hermitian regime
@@
Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model [30] and the PT-symmetric Aubry-André model [78, 38, 24, 25, 79, 44]. Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum [43, 78, 38, 25], topological phase transitions [44], and the localization characteristics of eigenstates [26, 63, 24, 79, 30, 91]. Given these insights, it becomes pertinent to investigate the effects of non-Hermiticity on our system.
@@
Expanding our setup into the non-Hermitian regime involves introducing spin-dependent dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam [91]. In such system, one generally focus on the post-selected space, which corresponds to the remaining atoms. Under this condition, both the eigenspectrum and the time evolution can be well descibed by the effective non-Hermtian Hamiltonian [37, 5]. Without loss of the generality, in the therotical simulation, we consider a system with balanced gain and loss term ($\gamma_{\uparrow}=-\gamma_{\downarrow}=\gamma$). We focus on the scenario with $M_{z,\uparrow}/M_{z,\downarrow}=-1$, which exhibits a critical phase in the Hermitian context. The mean fractal dimension for the system with non-Hermitian term ($\gamma=2$) is illustrated in Fig. 7(a). The region with $0<\bar{\eta}<1$ vanishes, potentially indicating that the critical phases in the Hermitian regime is suppressed. For system with larger $M_{z}$, the mean fractal dimension decreases to a near zero value, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises to near one value, signifying a transformation towards extended states. To avoid the finite size effect, we further present the fraction of extended, critical and localized phases for varying dissipation values in Fig. 7(b) with $t_{\text{so}}=0.5$. In the non-Hermitian regime ($\gamma\neq 0$), the critical phase is significantly suppressed, and the critical phase changed into a mixed phase with coexisting localized and extended states. Furthermore, we can also find as the non-Hermitian term increases, the pure localized and extended phases are also enlarged (Fig. 7 (c)).
The mechanism for the emergence of critical states in the Hermitian regime of this spinful model also provides insight for their suppression under non-Hermitian perturbations. As revealed in Refs.[93, 94], critical states in such systems arise when the on-site potential exhibits generalized incommensurate zeros in its $\sigma_{z}$ component. These zeros can be mapped to incommensurately distributed zeros in the hopping coefficients of an effective one-dimensional spinless model. The presence of such zeros effectively fragments the system into decoupled segments, forcing the delocalized wave functions to reorganize and develop the characteristic self-similar structure of critical states. However, the introduction of non-Hermitian terms disrupts the condition for generalized incommensurate zeros in the on-site potential. Consequently, the formation of critical states is invariably suppressed.
@@
Figure 7: Suppression of critical phase in non-Hermitian Raman lattice. (a) Mean fractal dimension $\bar{\eta}$ of the system under a non-Hermtian term ($\gamma=2$) when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. The dashed lines show the boundary between different phases in Hermitian regime. (b) Fraction of eigenstates belonging to the extended, critical, and localized phases as a function of $M_{z}$ for different values of $\gamma$ at $t_{\text{so}}=0.5$. In the non-Hermitian regime ($\gamma\neq 0$), the critical phase is significantly suppressed. (c) As the system enters the non-Hermitian regime, the critical phase evolves into a mixed phase with coexisting localized and extended states. The purely localized and extended phases are also enlarged with increasing non-Hermitian strength.
VII Conclusion and outlook
This study proposes a versatile Raman lattice configuration for theoretically investigating localization behavior for alkaline-earth-like atoms, using 173Yb as an example. Unlike previous approaches that primarily employed alkali-metal atoms [69, 70, 94], 173Yb represents a practically demonstrated fermionic system in optical Raman lattices [65, 64, 87]. The fermionic nature of 173Yb, along with the potential SU(N) symmetry of alkaline-earth atoms, can give rise to phenomena distinct from bosonic systems, particularly in the many-body regime [51, 68, 69].
@@ VII Conclusion and outlook
In this theoretical framework, we explore three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime. In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, our model predicts the emergence of a critical phase, which can be theoretically detected through signatures in expansion dynamics and spin polarization evolution. Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, theoretical calculations reveal a new phase with coexistence of extended, critical, and localized phases separated by mobility edges. Furthermore, we demonstrate that introducing non-Hermitian dissipation disrupts the generalized incommensurate zeros and further suppresses the critical phase. This work bridges a gap by providing a practical experimental scheme for implementing and studying localization with fermionic alkaline-earth atoms in optical Raman lattices, directly addressing experimental capabilities. The proposed scheme also facilitates exploration of the interplay between localization behavior and non-Hermiticity, thereby laying a foundation for future experimental studies with ultracold atoms [91, 30].
Acknowledgements.
@@
We thank Xin-Chi Zhou and Xiong-Jun Liu for fruitful discussions. GBJ acknowledges support from the RGC through 16302123, 16305024 and RFS2122-6S04.
Appendix A Finite size analysis
Figure 8: Finite size effects on $N_{e},N_{c},N_{l}$ and $\eta$. (a)-(c) Fractal dimension $\eta$ of individual eigenstates computed at three system sizes $L$, corresponding to $M_{z}=1.0,2.0,3.0$ respectively. (d)-(f) Scaling of $\eta$ with system size $L$ across different zones. (g)-(i) Fractions of extended, critical, and localized states ($N_{e}/N,N_{c}/N,N_{l}/N$) as functions of system size $L$. All results are simulated with $\gamma=1.0$ and $t_{\text{so}}=0.5$ in periodic boundary conditions.
In this section, we investigate the finite size effects on the fractal dimension and the fractions of extended, critical, and localized states ($N_{e}/N,N_{c}/N,N_{l}/N$) in the non-Hermitian regime. Figure 8(a)-(c) shows the fractal dimension $\eta$ for $t_{\text{so}}=0.5$ and $\gamma=1.0$ at different system sizes $L$. For lower $M_{z}$ (e.g., $M_{z}=1.0$), the fractal dimensions of all eigenstates increase with system size and approach 1, indicating extended behavior. Conversely, for higher $M_{z}$ (e.g., $M_{z}=3.0$), the fractal dimensions decrease with increasing $L$ and tend to zero, consistent with behaviors of localized states. In the intermediate regime ($M_{z}=2.0$), eigenstates with larger absolute eigenenergies ($|E|>1.15$) exhibit decreasing fractal dimensions as $L$ increases, while those with lower energies show the opposite trend, suggesting a phase with coexisting extended and localized states.
Based on these observations, we classify eigenstates into distinct zones sharing similar localization properties. The average fractal dimension within each zone is presented in Figure 8(d)-(f). For $M_{z}=1.0$ and $3.0$, all eigenstates fall into a single zone, corresponding to a pure phase. At $M_{z}=2.0$, however, eigenstates separate into two zones with mean fractal dimensions evolving oppositely with system size—one increasing and the other decreasing. This allows us to assign consistent localization properties to all states within each zone and determine the fractions of extended, critical, and localized states. As shown in Figure 8(g)-(i), these fractions remain invariant with increasing system size.
Appendix B Absence of non-Hermitian skin effect
In this section, we validate the absence of the non-Hermitian skin effect in our system, confirming that the observed suppression of the critical phase does not originate from the skin effect. According to the established correspondence between winding numbers and skin modes in non-Hermitian systems [81, 87, 54, 90], a system exhibits the skin effect under open boundary conditions (OBC) if and only if the eigenvalue winding on the complex plane displays nontrivial closed-loop topology under periodic boundary conditions (PBC). The presence of the skin effect also leads to distinct eigenspectra under PBC and OBC. As shown in Figure 9, the eigenenergy spectra on the complex plane are presented for both boundary conditions. The PBC spectrum exhibits no loop structures, indicating the absence of the non-Hermitian skin effect in the present system.
Figure 9: Eigenenergy spectrum on complex plane under PBC and OBC with $M_{z}=1.0,2.0,3.0$ respectively. All results are simulated with $\gamma=1.0$ and $t_{\text{so}}=0.5$.
Appendix C Effect of the width of the initial wave packet on the expansion dynamics
Figure 10: Effect of the width of initial wave packet. (a-d) Expansion dynamics of the wave packet, characterized by $W$ as a function of $M_{z}$, with $t_{so}=0.3$ and an initial width of $a=5,10,20,40$, respectively. (e) $W$ as a function of time $t$ for different widths with a fixed $M_{z}=0.5$. Other parameters are the same as the main Fig. 3.
Figure 11: Effect of the total evolution time. (a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}$ = 0.5), critical phase ($M_{z}$ = 2.0), and localized phase ($M_{z}$ = 3.5) with the initial state comprising spin-up atoms in a uniform distribution.(b) Time-averaged spin polarization $\bar{I}_{p}$ as a function of $M_{z}$ averaged over different time interval $t_{f}=10,30,100,200$.(c) Time-averaged density imbalance $\bar{I}_{d}$ as a function of $M_{z}$ averaged over different time interval $t_{f}=10,30,100,200$. Other parameters are identical to those in Fig.5(d).
In this section, we investigate the effect of the initial wave packet width $a$ on the expansion dynamics. We find that while the initial wave packet width $a$ does not alter the qualitative distinction in dynamical behaviors among different phases, it does quantitatively influence the clarity of the contrast and the expansion speed, as shown in Figure 10.
Appendix D Effect of the evolution time on time average $\bar{I}_{d}$ and $\bar{I}_{p}$
In this section, we examine the influence of the total evolution time on the dynamics of the density imbalance and spin polarization. We observe that while a very short evolution time can quantitatively affect the time-averaged spin polarization and density imbalance, the qualitative trends across different phases remain unchanged, as demonstrated in Figure 11. Moreover, the time-averaged spin polarization is robust against the total evolution time once it spans several oscillation periods. Hence, despite experimental constraints that may limit the coherent evolution time, spin dynamics can still serve as a reliable detection scheme for distinguishing between phases—provided the total evolution time is sufficiently long to encompass several periods of spin oscillation.